OpenStudy (anonymous):

Please explain how to solve this. Find an for each arithmetic sequence: a1=5; d=4; n=1

hartnn (hartnn):

general formula for an of arithmetic sequence is \(\large a_n-a_1+(n-1)d\) just plug in values!

hartnn (hartnn):

its, \(\large a_n=a_1+(n-1)d\) with a1= 5, d= 4

hartnn (hartnn):

did you try to put a1=5,d=4 in that formula?

OpenStudy (anonymous):

is the answer 9? i'm not really sure...

hartnn (hartnn):

9 for what ? you need to find general tern \(a_n\) right ? it'll be in terms of 'n' \(\large a_n=a_1+(n-1)d \\ with \: \: a_1=5,d=4, \\ a_n=5+(n-1)4\) can you simplify that ?

OpenStudy (anonymous):

i'm supposed to find \[a _{n}\]

hartnn (hartnn):

yes, and thats what i wrote the formula for.. \(a_n=a_1+(n-1)d\)

hartnn (hartnn):

didn't you get this ? \(\large a_n=a_1+(n-1)d \\ with \: \: a_1=5,d=4, \\ a_n=5+(n-1)4\)

OpenStudy (anonymous):

yes, but shouldn't i substitute n with 1? is the answer 5?

hartnn (hartnn):

if you want to find \(a_1\) from \(a_n\), then you plug in n=1. but thats not required, a1 is already given to be 5 you need to find \(a_n\) which you'll get after simplifying \(a_n=5+(n-1)4\)

OpenStudy (anonymous):

oh, ok...